# & Télécom Paris, C2

### v5

@@ 5 The distinguisher
In [5, §3.4], the security level $2^{b}$ — or less formally, the number $b$ of targeted “bits of security” — of the Classic McEliece system, essentially defined as the complexity of the currently best attack (namely, by information set decoding), is shown asymptotically to be linear in $t\propto n/\log(n)$. Then in (92) we have
$\frac{\log(\kappa)}{n/\log(n)}\propto\frac{(\log\log(n))^{3}}{\log(n)}\to 0$
6 Conclusion and open problems
We presented the first structural analysis of the McEliece cryptosystem whose asymptotic complexity is better than that of generic decoding algorithms; more precisely, subexponential in the error-correcting capability of the code. However this is only an asymptotic result: for concrete, finite parameters, such as those proposed in the Classic McEliece specification, a naive implementation of our distinguisher still falls beyond the best attacks by a non-negligible factor.
Problem 1
Improve the implementation of this distinguisher.333addendum: actually, the improvements discussed below can already be found in works of Kreuzer-Lusteck and Albano-La Scala
For instance, if we choose a monomial order in Algorithm 1, and have a reduced basis for one syzygy space, then the generalized Macaulay matrix constructed from it will be (i) somewhat sparse and (ii) already partially reduced. This could be exploited to make the computation of (a reduced basis of) the next syzygy space faster.
Alternatively, instead of iteratively computing a whole range of the linear strand, one could directly compute a targeted Betti number, e.g. using Koszul cohomology (the author would like to thank W. Castryck for this suggestion). It is then likely that this Koszul cohomology computation could also benefit from similar optimizations (i) and (ii) exploiting sparsity and structure.
Problem 2
Can our distinguisher, and in particular the Betti number computation, benefit from a quantum speedup?
Problem 3
Our lower bounds on the Betti numbers, and in particular the $r_{\max}$, of dual alternant or Goppa codes, are not tight in general. Can they be improved? Better, can one give a complete, explicit description of the minimal resolution of these codes? (See at the beginning of the Supplementary material for elements in this direction.)
We initiated the study of higher syzygies of codes (or equivalently, of finite sets of points in projective space) from a genuinely coding theoretical perspective.
Problem 4
Pursue this study in a systematic way. This should cover basic operations on codes (some of which have a nice geometric interpretation, some of which less so), and also crucially, make links with various metric properties of the code such as its weight distribution, generalized weight hierarchy, etc.
Problem 5
Can higher syzygies serve in decoding algorithms?
Problem 6
Betti numbers actually are invariants of the monomial equivalence class of a code (Lemma 3). Can they find other applications, for instance, to problems of code equivalence?
Problem 7
Can our distinguisher be turned into a key recovery attack, or more properly, can some of the ideas beneath be used to build a key recovery attack? For instance, one would like to combine our approach with:
  • filtration arguments
  • extraction of short relations, e.g. via MinRank techniques.
(Observe that some of the explicit syzygies described in Proposition 3 are short, while Proposition 4 shows that short relations pass to shortened subcodes.)
It is very common that new mathematical tools (Euclidean lattices, elliptic curves, pairings, isogenies…) are introduced in the cryptographic realm first for cryptanalytic purposes, i.e. to break systems. Once digested by the community, they are then used in a more constructive way, to build new cryptosystems.
Problem 8
Can one devise syzygy-based cryptography?
{credits}
6.0.1 Acknowledgements
Supplementary material
More on syzygies of shortened dual alternant and Goppa codes
It is possible to improve the lower bound on $\beta_{r-1,r}({\mathsf{C}}_{s})$ in Theorem 3.1, using not only $\mathbf{\Phi}=\mathbf{\Phi}^{(0)}$, but also its conjugates $\mathbf{\Phi}^{(1)},\dots,\mathbf{\Phi}^{(m-1)}$.
Let us first consider the simpler case $s=0$.
The Eagon-Northcott complex of each such conjugate gives a subcomplex of the minimal resolution of ${\mathsf{C}}$, however one should take care of the fact that in general these subcomplexes are not in direct sum. Proposition 3 gives a basis $s^{(j)}_{r;i_{1},\dots,i_{r}}(\mathbf{\Phi}^{(u)})$ for each syzygy space $M_{r-1,r}(\mathbf{\Phi}^{(u)})$. The indices $i_{1},\dots,i_{r}$ correspond to columns of $\mathbf{\Phi}^{(u)}=\left(\begin{array}[]{c|c|c|c}\mathbf{B}^{(u+e)}_{0}&\mathbf{B}^{(u+e-1)}_{1}&\cdots&\mathbf{B}^{(u)}_{e}\end{array}\right)$. It is then easily seen:
Proposition 9
The $s^{(j)}_{r;i_{1},\dots,i_{r}}(\mathbf{\Phi}^{(u)})$ with column $i_{1}$ not in $\mathbf{B}^{(e+u)}_{0}$ also belong to $M_{r-1,r}(\mathbf{\Phi}^{(u-1)})$. Conversely, for $e<m/2$, those with $i_{1}$ in $\mathbf{B}^{(e+u)}_{0}$ are linearly independent, hence form a basis of the subspace $V_{r}=M_{r-1,r}(\mathbf{\Phi}^{(0)})+\cdots+M_{r-1,r}(\mathbf{\Phi}^{(r)})$ of $M_{r-1,r}({\mathsf{C}})$.
This gives the improved lower bound:
Corollary 6
Set $e=\lfloor\log_{q}(t-1)\rfloor$ and $f=(e+1)t-\frac{q^{e+1}-1}{q-1}$. Then for any $r\geq 2$ we have
$\beta_{r-1,r}({\mathsf{C}})\;\geq\;\dim(V_{r})=m(r-1)\left(\binom{f}{r}-\binom{f-(t-1)}{r}\right).$ (97)
For $r=2$ this specializes to
$\beta_{1,2}({\mathsf{C}})\;\geq\;\dim(V_{2})=\frac{m(t-1)}{2}\left((2e+1)t-2\frac{q^{e+1}-1}{q-1}\right),$ (98)
a result already proved in [12] by (essentially) the same method. It is observed moreover than in many cases the lower bound (98) is an equality, so we actually have
$I_{2}({\mathsf{C}})=V_{2}.$ (99)
Likewise for $r>f-(t-1)$ this gives
$\beta_{r-1,r}({\mathsf{C}})\;\geq\;\dim(V_{f})=m(r-1)\binom{f}{r}$ (100)
and we will see examples where this is an equality.
However for arbitrary $r$, the improved lower bound (97) still isn’t tight, because the inclusion $V_{r}\subseteq M_{r-1,r}({\mathsf{C}})$ is strict in general. For instance, one can show that $M_{2,3}({\mathsf{C}})$ contains syzygies of the form
$\begin{split}X_{a}(&X_{b+1}^{(u)}X_{c+1}^{(u+v)}-X_{b+q^{v}+1}^{(u)}X_{c}^{(u+v)})-X_{a+q^{u}}(X_{b}^{(u)}X_{c+1}^{(u+v)}-X_{b+q^{v}}^{(u)}X_{c}^{(u+v)})+\\ &+X_{c}^{(u+v)}(X_{a}X_{b+q^{v}+1}^{(u)}-X_{a+q^{u}}X_{b+q^{v}}^{(u)})-X_{c+1}^{(u+v)}(X_{a}X_{b+1}^{(u)}-X_{a+q^{u}}X_{b}^{(u)})=0\end{split}$ (101)
that do not belong to $V_{3}$ in general.
In principle, under (99) it should be possible to use the explicit description of $V_{2}$ together with Proposition 2 to compute (the linear strand of) the minimal resolution of ${\mathsf{C}}$. However the combinatorics appears quite complicated.
Now for general $s$, (68) and (73) can be extended to:
Experimental fact 3
Let ${\mathsf{C}}$ be a dual alternant code. Then for all $s$ we have
$r_{\max}({\mathsf{C}}_{s})\geq r_{\max}({\mathsf{C}})-s.$ (102)
It would be tempting to conjecture that (102) holds for all codes, but it turns out that one can find counterexamples. However, these counterexamples are quite rare. So maybe an interesting problem instead should be to give criteria for (102) to hold.
Observe that shortening is dual to puncturing. As the behaviour of minimal resolutions of codes under puncturing is quite well understood (Corollary 3), maybe the study of shortening should go hand-in-hand with that of code duality. Experiments suggest a loose link between $\beta_{r-2,r}({\mathsf{C}}^{\perp})$ and $\beta_{k-r,k+1-r}(C)$, at least for some regimes of parameters.
Back to dual alternant or Goppa codes, we actually observe a strong regular pattern (for Goppa codes the author only tested the irreducible case, but the result is likely to generalize):
Experimental fact 4
Let $\mathcal{T}=\operatorname{{\mathsf{Alt}}}^{\perp}$ or $\operatorname{{\mathsf{Gop}}}^{\operatorname{irr},\perp}$ be a type of codes, namely, either dual alternant codes, or dual Goppa codes with irreducible Goppa polynomial. Let $q$ be a field cardinality, and $t\geq 3$ an integer.
  1. 1.
    For all $m$ large enough,
    $r_{\max}({\mathsf{C}})=r^{*}_{\mathcal{T},q,t}$ (103)
    is the same for generic proper ${\mathsf{C}}\in\mathcal{T}_{q,m,q^{m},t}$, i.e. it generically does not depend on $m$ nor on the choice of ${\mathsf{C}}$, but only on $\mathcal{T},q,t$.
Now $\mathcal{T},q,t$ being fixed, we set $r^{*}=r^{*}_{\mathcal{T},q,t}$. Also, given $m,t$, we set $k=mt$.
  1. 2.
    For all $m$, for all $n\leq q^{m}$, for generic proper ${\mathsf{C}}\in\mathcal{T}_{q,m,n,t}$, and for all $s\leq r^{*}-2$, if $r^{*}-s>\max\left(\frac{(k-s)(k-s+1)}{n-s},k-s+1-\operatorname{d}_{\min}({\mathsf{C}}_{s})\right)$, then
    $r_{\max}({\mathsf{C}}_{s})=r^{*}-s.$ (104)
  2. 3.
    For $20\leq i\leq r^{*}-$ 2 there are functions $b_{i}(r)$ (actually depending on $\mathcal{T},q,t$, so $b_{i}(r)=b_{i,\mathcal{T},q,t}(r)$) such that, for all $m$, for all $n\leq q^{m}$, for generic proper ${\mathsf{C}}\in\mathcal{T}_{q,m,n,t}$, for all $s\leq r^{*}-2$, and for all $r$ in the interval $2\leq r\leq r^{*}-s$, if $r>\max\left(\frac{(k-s)(k-s+1)}{n-s},k-s+1-\operatorname{d}_{\min}({\mathsf{C}}_{s})\right)$, then
    $\beta_{r-1,r}({\mathsf{C}}_{s})=m\,b_{r^{*}-s-r}(r).$ (105)
  3. 4.
    We have $b_{0}(r)=r-1$ independently of $\mathcal{T},q,t$, so (105) reduces to
    $\beta_{r^{*}-s-1,r^{*}-s}({\mathsf{C}}_{s})=m\,(r^{*}-s-1).$ (106)
    More generally for $i\geq 1$ we have $b_{i}(r)=(r-1)\binom{r+i}{i}$ for a certain number of (but now, not for all) values of $\mathcal{T},q,t$, so then (105) reduces to
    $\beta_{r-1,r}({\mathsf{C}}_{s})=m\,(r-1)\binom{r^{*}-s}{r}.$ (107)
We observe that (68) gives a lower bound on $r^{*}$. For alternant codes, experimentally, this lower bound is an equality. However for Goppa codes this lower bound seems to be always a strict inequality. And even for binary Goppa codes, the improved lower bound (73) still is a strict inequality.
$\begin{array}[]{c|ccccccc}s&\beta_{1,2}&\beta_{2,3}&\beta_{3,4}&\beta_{4,5}&\beta_{5,6}&\beta_{6,7}&\beta_{7,8}\\ \hline\cr 0&251&1400&3230&2480&\mathbf{1400}&\mathbf{480}&\mathbf{70}\\ 1&202&880&1170&\mathbf{840}&\mathbf{350}&\mathbf{60}&-\\ 2&154&440&\mathbf{450}&\mathbf{240}&\mathbf{50}&-&-\\ 3&107&\mathbf{200}&\mathbf{150}&\mathbf{40}&-&-&-\\ 4&66&\mathbf{80}&\mathbf{30}&-&-&-&-\\ 5&31&\mathbf{20}&-&-&-&-&-\\ 6&\mathbf{10}&-&-&-&-&-&-\\ 7&-&-&-&-&-&-&-\end{array}$
Figure 4: $s$-shortened $\operatorname{{\mathsf{Alt}}}^{\perp}_{2,10,5}\;$ ($r^{*}=8$)
$\begin{array}[]{c|ccccccc}s&\beta_{1,2}&\beta_{2,3}&\beta_{3,4}&\beta_{4,5}&\beta_{5,6}&\beta_{6,7}&\beta_{7,8}\\ \hline\cr 0&222&1943&1725&\mathbf{1120}&\mathbf{700}&\mathbf{240}&\mathbf{35}\\ 1&193&1344&\mathbf{525}&\mathbf{420}&\mathbf{175}&\mathbf{30}&-\\ 2&165&801&\mathbf{225}&\mathbf{120}&\mathbf{25}&-&-\\ 3&138&312&\mathbf{75}&\mathbf{20}&-&-&-\\ 4&112&\mathbf{40}&\mathbf{15}&-&-&-&-\\ 5&87&\mathbf{10}&-&-&-&-&-\\ 6&63&-&-&-&-&-&-\\ 7&40&-&-&-&-&-&-\\ \end{array}$
Figure 5: $s$-shortened $\operatorname{{\mathsf{Alt}}}^{\perp}_{3,5,6}\;$ ($r^{*}=8$)
$\begin{array}[]{c|ccccccccc}s&\beta_{1,2}&\beta_{2,3}&\beta_{3,4}&\beta_{4,5}&\beta_{5,6}&\beta_{6,7}&\beta_{7,8}&\beta_{8,9}\\ \hline\cr\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 8&280&3224&7464&4272&\mathbf{3360}&\mathbf{1728}&\mathbf{504}&\mathbf{64}\\ 9&249&2510&1800&\mathbf{1792}&\mathbf{1120}&\mathbf{384}&\mathbf{56}&-\\ 10&219&1856&\mathbf{840}&\mathbf{672}&\mathbf{280}&\mathbf{48}&-&-\\ 11&190&1260&\mathbf{360}&\mathbf{192}&\mathbf{40}&-&-&-\\ 12&162&720&\mathbf{120}&\mathbf{32}&-&-&-&-\\ 13&135&234&\mathbf{24}&-&-&-&-&-\\ 14&109&\mathbf{16}&-&-&-&-&-&-\end{array}$
Figure 6: $s$-shortened $\operatorname{{\mathsf{Gop}}}^{\perp}_{2,8,5}\;$ ($r^{*}=17$)
$\begin{array}[]{c|cccccc}s&\beta_{1,2}&\beta_{2,3}&\beta_{3,4}&\beta_{4,5}&\beta_{5,6}&\beta_{6,7}\\ \hline\cr 0&180&1293&3090&3144&960&\mathbf{36}\\ 1&151&810&1350&732&\mathbf{30}&-\\ 2&123&450&450&\mathbf{24}&-&-\\ 3&96&210&24&-&-&-\\ 4&70&30&-&-&-&-\\ 5&45&-&-&-&-&-\\ 6&21&-&-&-&-&-\\ 7&-&-&-&-&-&-\end{array}$
Figure 7: $s$-shortened $\operatorname{{\mathsf{Gop}}}^{\perp}_{3,6,5}\;$ ($r^{*}=7$)
$\begin{array}[]{c|cccccc}s&\beta_{1,2}&\beta_{2,3}&\beta_{3,4}&\beta_{4,5}&\beta_{5,6}&\beta_{6,7}\\ \hline\cr\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 32&719&8474&\mathbf{450}&\mathbf{240}&\mathbf{50}&-\\ 33&662&6216&\mathbf{150}&\mathbf{40}&-&-\\ 34&606&4070&\mathbf{30}&-&-&-\\ 35&551&2034&-&-&-&-\\ \end{array}$
Figure 8: $s$-shortened $\operatorname{{\mathsf{Gop}}}^{\perp}_{2,10,9}\;$ ($r^{*}=38$)
Figures 8-8 illustrate the last Experimental fact. Boldface values match (107). This strongly suggests that the minimal resolution of the codes contains $m$ conjugate Eagon-Northcott complexes of length $r^{*}$, whose top degree components are in direct sum.
Computing the Betti diagram of some families of codes
Lemma 12 ([15, Th. 1], reformulated)
Proposition 10
$\beta_{r-1,r}=\frac{(r-1)(k-r)}{k}\binom{k+1}{r}$ (108)
for $12\leq r\leq k-$ 1. In particular they satisfy the symmetry $\beta_{i,j}=\beta_{k-1-i,k+1-j}$.
Proof
Proposition 11
Let ${\mathsf{C}}$ be a $[2k-1,k]$ $\operatorname{{\mathsf{GRS}}}$ code. Then the nonzero Betti numbers of ${\mathsf{C}}$ are:
  • $\beta_{0,0}=1$
  • $\beta_{r-1,r}=(r-1)\binom{k-1}{r}\;$ for $12\leq r\leq k-$ 1
Proof
Now set $n=2k-1$.
Figures 10 and 10 illustrate Propositions 10 and 11.
Figure 9: a $[9,8]$ parity or $\operatorname{{\mathsf{GRS}}}$ code
Figure 10: a $[15,8]$ $\operatorname{{\mathsf{GRS}}}$ code
Exercise 1
Compute the Betti diagram of $[n,k]$ $\operatorname{{\mathsf{GRS}}}$ codes for $k+1<n<2k-1$.
Experimental data on defects for random codes
Figures 13-13 compare Betti diagrams for various types of codes of length $n=23$ and dimension $k=12$. Observe that these parameters give $\frac{k(k+1)}{n}\approx 6.78$.
First, Figure 13 presents the Betti diagram as predicted by the minimal resolution conjecture, i.e. with all defects equal to zero. It turns out that the minimal resolution conjecture is true for these parameters, so generic $[23,12]$-codes over an infinite field actually have this Betti diagram. Such codes are MDS, i.e. they have $d=12$ and $d^{\perp}=13$, so the intervals $d^{\perp}\leq r\leq\frac{k(k+1)}{n}$ and $\frac{k(k+1)}{n}\leq\ r\leq k+1-d$ are empty, and Experimental fact 1 is trivially verified. Actually, codes with this Betti diagram can be found even over not-so-large finite fields, e.g. $q=5$ suffices (the author did not search for optimality). These codes are no longer MDS, but they still have $d,d^{\perp}\geq 7$ and again Experimental fact 1 is trivially verified. Moreover, the converse Experimental fact 2 is also verified as all defects are zero.
$\begin{array}[]{c|cccccccccccc}&0&1&2&3&4&5&6&7&8&9&10&11\\ \hline\cr 0&1&-&-&-&-&-&-&-&-&-&-&-\\ 1&-&55&319&880&1353&990&-&-&-&-&-&-\\ 2&-&-&-&-&-&330&1617&1870&1221&485&110&11\\ \end{array}$
@@ 6.0.1 Acknowledgements
$\begin{array}[]{c|cccccccccccc}&0&1&2&3&4&5&6&7&8&9&10&11\\ \hline\cr 0&1&-&-&-&-&-&-&-&-&-&-&-\\ 1&-&55&319&884&1397&1224&490&121&18&1&-&-\\ 2&-&-&4&44&234&820&1738&1888&1222&485&110&11\\ \end{array}$
Figure 12: a (pseudo)random $[23,12]_{2}$-code ($d=3$, $d^{\perp}=4$)
$\begin{array}[]{c|cccccccccccc}&0&1&2&3&4&5&6&7&8&9&10&11\\ \hline\cr 0&1&-&-&-&-&-&-&-&-&-&-&-\\ 1&-&55&320&891&1408&1210&320&55&-&-&-&-\\ 2&-&1&11&55&220&650&1672&1870&1221&485&110&11\\ \end{array}$
Figure 13: the $[23,12]_{2}$ Golay code ($d=7$, $d^{\perp}=8$)
@@
More systematically, Figure 14 presents statistics on $\operatorname{def}(\varphi_{r})$ ($2\leq r\leq 8$) for random $[56,16]_{2}$-codes. For each pair $(d,d^{\perp})$, a few thousands of codes with these parameters were sampled uniformly (using rejection sampling). The average value of $\operatorname{def}(\varphi_{r})$ among these samples is displayed, and also its $99\%$ distribution interval (which means at most $0.5\%$ fall below and at most $0.5\%$ above).
Here we have $\frac{k(k+1)}{n}\approx 4.86$. Then, in accordance with Experimental fact 1, we can check $\operatorname{def}(\varphi_{r})>0$ for $d^{\perp}\leq r\leq 4$ and for $5\leq r\leq 17-d$ (when applicable); while $\operatorname{def}(\varphi_{r})=0$ with higher and higher probability as we move away from these intervals, in accordance with Experimental fact 2.
@@
Figure 14: some experimental data on $\operatorname{def}(\varphi_{r})$ for random $[56,16]_{2}$-codes
@@
We experimentally find that dual Goppa codes ${\mathsf{C}}$ with these parameters, with $n=q^{m}=256$, consistently have $r_{\max}=4$, which is better than the lower bound in Theorem 3.1 ($e=0$, $r_{\max}\geq t-1=3$). Moreover, their Betti numbers are:
@@
$\begin{array}[]{|c|ccccccccccccc|}\hline\cr n&\cdots&88&87&\mathbf{86}&85&84&\cdots&70&69&\mathbf{68}&67&66&\cdots\\ \hline\cr\beta_{2,3}^{\operatorname{{\mathsf{Gop}}}^{\text{irr},\perp}}&&80&80&\mathbf{80}&85&100&&310&325&340&355&370&\\ \beta_{2,3}^{\text{random}}&&40&55&\mathbf{70}&85&100&&310&325&340&355&370&\\ \hline\cr\beta_{3,4}^{\operatorname{{\mathsf{Gop}}}^{\text{irr},\perp}}&&12&12&12&12&12&&12&12&\mathbf{12}&105&210&\\ \beta_{3,4}^{\text{random}}&&0&0&0&0&0&&0&0&\mathbf{0}&105&210&\\ \hline\cr\end{array}$
@@
We experimentally find that dual Goppa codes ${\mathsf{C}}$ with these parameters, with $n=q^{m}=64$, consistently have $r_{\max}=8$, which is better than the lower bound in Corollary 4 ($\widehat{e}=1$, $r_{\max}\geq 7$). Moreover, their top Betti numbers are:
@@
What happens? It turns out the conditions in Heuristic 1 are not satisfied anymore, so we cannot ensure $\operatorname{def}(\varphi_{r})=0$, or even $\operatorname{def}(\varphi_{r})$ small, for these values of $r$ and $n$. Indeed, in order to have $\beta_{r-1,r}=0$, we need $r>k+1-d$, where $d=\operatorname{d}_{\min}({\mathsf{C}})$. For $r=7$ and $k=mt=18$ this gives $d\geq 13$. Then for smaller $d$, from Experimental fact 1, we expect a loose link between $\beta_{r-1,r}$ and the weight distribution of ${\mathsf{C}}$. Recall the distinguisher still works as long as random codes satisfy $\beta_{r-1,r}<\beta_{r-1,r}^{*}$ with high enough probablity. Experimentally, for $r=7$, we find this inequality is satisfied for random codes of minimum distance $d\geq 10$, while $d=9$ is a borderline case, and $d\leq 8$ fails invariably. Now what happens is that, as $n$ decreases from $56$ to $50$, the proportion of random codes with $d\geq 10$ also decreases, and they become minority for $n=49$.
@@
The sheer fact that the minimum distance is easily computable for codes with such small parameters actually provides a much more efficient alternative distinguisher. Indeed, turning back to the primal codes, the designed distance for binary Goppa codes with $t=3$ is $7$. Let us thus consider a distinguisher that decides that a given $[n,n-18]$-code is Goppa if it has $\operatorname{d}_{\min}\geq 7$, and that it is random otherwise. Experiments then show that, as long as $n\geq 32$, less than $2\%$ of random codes with these parameters have $\operatorname{d}_{\min}\geq 7$. This means that this distinguisher has more than $99\%$ success rate! And it remains remarkably effective even for smaller $n$. For instance, as low as $n=24$, less than half random codes have $\operatorname{d}_{\min}\geq 7$, so this distinguisher still has more than $75\%$ success rate.

### v6

@@ 4 Regularity $2$ and the small defect heuristic
[…]
Figure 12: a (pseudo)random $23,12]_{2}-$code ($d=3$, $d^{\perp}=4$)
[…]
Figure 13: the $23,12]_{2}$ Golay code ($d=7$, $d^{\perp}=8$)
@@
More systematically, Figure 14 presents statistics on $\operatorname{def}(\varphi_{r})$ ($2\leq r\leq 8$) for random $56,16]_{2}-$codes. For each pair $(d,d^{\perp})$, a few thousands of codes with these parameters were sampled uniformly (using rejection sampling). The average value of $\operatorname{def}(\varphi_{r})$ among these samples is displayed, and also its $99\%$ distribution interval (which means at most $0.5\%$ fall below and at most $0.5\%$ above).
Here we have $r_{\operatorname{crit}}\approx 4.86$. Then, in accordance with Experimental fact 1, we can check $\operatorname{def}(\varphi_{r})>0$ for $d^{\perp}\leq r\leq 4$ and for $5\leq r\leq 17-d$ (when applicable); while $\operatorname{def}(\varphi_{r})=0$ with higher and higher probability as we move away from these intervals, in accordance with Experimental fact 2.
Experiments with other parameters display similar results.
@@
Figure 14: some experimental data on $\operatorname{def}(\varphi_{r})$ for random $56,16]_{2}-$codes
Consequently, we postulate the following relaxation of the minimal resolution conjecture:
Heuristic 1
Fix a field cardinality $q$, assume $n$ is not too close to $k$ in order to stay away from the counterexamples to the minimal resolution conjecture (e.g. $n>k+\sqrt{2k}+2$ should suffice), and $n<\binom{k+1}{2}$ in order to ensure regularity $2$. Then for random $n,k]_{q}-$codes, with high probability:
  1. 1.
    if $d>k+1-r_{\operatorname{crit}}$ we expect $\beta_{r-1,r}=0$ for $r>r_{\operatorname{crit}}$
  2. 2.
    if $d^{\perp}>r_{\operatorname{crit}}$ we expect $\beta_{r-1,r}=\left(\frac{k(k+1)}{r}-n\right)\binom{k-1}{r-2}$ for $r<r_{\operatorname{crit}}$
where $r_{\operatorname{crit}}=\frac{k(k+1)}{n}$.
Remark 3
5 The distinguisher
Basic version (without shortening)
In the following paragraphs we compare our new distinguisher (in this basic version) with the one from [10], on the very same examples taken from this reference: $\operatorname{{\mathsf{Gop}}}^{\text{irr}}_{4,4,4}$ and $\operatorname{{\mathsf{Gop}}}^{\text{irr}}_{2,6,3}$. In both cases, we improve the distinguishability threshold. Interestingly, for $\operatorname{{\mathsf{Gop}}}^{\text{irr}}_{4,4,4}$, this improved threshold is the one given by (111) (with $r=r_{\max}=4$), i.e. the conditions in Heuristic 1 are always met. On the other hand, for $\operatorname{{\mathsf{Gop}}}^{\text{irr}}_{2,6,3}$, the conditions in Heuristic 1 are the limiting factor (but still improving over [10]).
@@ 5 The distinguisher
We experimentally find that dual Goppa codes $\mathsf{C}$} with these parameters, with $n=q^{m}=256$, consistently have $r_{\max}=4$, matching the nonbinary case at the end of Remark 1 ($a=\lceil\log_{q}(t/(q-1))\rceil=1$, $r^{*}=(a+1)t-q^{a}=4$). Moreover, their Betti numbers are:
@@
[…]
@@
We experimentally find that dual Goppa codes $\mathsf{C}$} with these parameters, with $n=q^{m}=64$, consistently have $r_{\max}=8$, matching the value given in Remark 1 ($a=\lceil\log_{2}(t))\rceil=2$, $r^{*}=(a+2)t-2^{a}=8$). Moreover, their top Betti numbers are:
@@
What happens? It turns out the conditions in Heuristic 1 are not satisfied anymore, so we cannot ensure $\operatorname{def}(\varphi_{r})=0$, or even $\operatorname{def}(\varphi_{r})$ small, for these values of $r$ and $n$. Indeed, in order to have $\beta_{r-1,r}=0$, we need r>k+1dr>k+1-d, where $d=\operatorname{d}_{\min}({\mathsf{C}})$. For $r=7$ and $k=mt=18$ this gives $d\geq 13$. Then for smaller $d$, from Experimental fact 1, we expect a loose link between $\beta_{r-1,r}$ and the weight distribution of $\mathsf{C}$}. Recall the distinguisher still works as long as random codes satisfy $\beta_{r-1,r}<\beta_{r-1,r}^{*}$ with high enough probablity. Experimentally, for $r=7$, we find this inequality is satisfied for random codes of minimum distance $d\geq 10$, while $d=9$ is a borderline case, and $d\leq 8$ fails invariably. Now what happens is that, as $n$ decreases from $56$ to $50$, the proportion of random codes with $d\geq 10$ also decreases, and they become minority for $n=49$.
@@
The sheer fact that the minimum distance is easily computable for codes with such small parameters actually provides a much more efficient alternative distinguisher. Indeed, turning back to the primal codes, the designed distance for binary Goppa codes with $t=3$ is $7$. Let us thus consider a distinguisher that decides that a given [n,n18][n,n-18]-code is Goppa if it has $\operatorname{d}_{\min}\geq 7$, and that it is random otherwise. Experiments then show that, as long as $n\geq 32$, less than $2\%$ of random codes with these parameters have $\operatorname{d}_{\min}\geq 7$. This means that this distinguisher has more than $99\%$ success rate! And it remains remarkably effective even for smaller $n$. For instance, as low as $n=24$, less than half random codes have $\operatorname{d}_{\min}\geq 7$, so this distinguisher still has more than $75\%$ success rate.